# What should a neuron aim for? Designing local objective functions based on information theory

### v3

@@ Abstract
In modern deep neural networks, the learning dynamics of the individual neurons is often obscure, as the networks are trained via global optimization. Conversely, biological systems build on self-organized, local learning, achieving robustness and efficiency with limited global information. We here show how self-organization between individual artificial neurons can be achieved by designing abstract bio-inspired local learning goals. These goals are parameterized using a recent extension of information theory, Partial Information Decomposition (PID), which decomposes the information that a set of information sources holds about an outcome into unique, redundant and synergistic contributions. Our framework enables neurons to locally shape the integration of information from various input classes, i.e. feedforward, feedback, and lateral, by selecting which of the three inputs should contribute uniquely, redundantly or synergistically to the output. This selection is expressed as a weighted sum of PID terms, which, for a given problem, can be directly derived from intuitive reasoning or via numerical optimization, offering a window into understanding task-relevant local information processing. Achieving neuron-level interpretability while enabling strong performance using local learning, our work advances a principled information-theoretic foundation for local learning strategies.
@@ 1 Introduction
Neuroscience models have indeed shown that local learning rules [1, 2, 3] can solve a variety of tasks, such as sequence learning, unsupervised clustering, and classification tasks [1, 4, 5]. Being often formulated on the basis of temporal coincidence (spike-timing), these very mechanistic rules do not offer direct insights into the actual information processing at each neuron. To facilitate such functional insight, more recently information theory has been used to formulate more abstract, but interpretable frameworks for local learning rules (e.g., Kay 6, Wibral et al. 7).
@@
Information theory [8] allows one to abstractly prescribe how much of the information from different inputs should be conveyed to the output of a neuron [6]. Moreover, it provides a general and flexible mathematical framework, which captures the information processing at an abstract and interpretable level free of details of the exact implementation. As such, information theory has been employed to define general optimization goals, including global objectives like cross-entropy loss [9] and more localized ones such as local greedy infomax [10], and early attempts at passing only information that is coherent between multiple inputs [11].
@@
Nevertheless, classical information theory falls short of describing how multiple sources work together to produce an output. In fact, it completely lacks the capacity to describe how the information from different sources is integrated in redundant, synergistic or unique ways in the output. These different ways, or information atoms, however, contribute very differently to a neuron’s function: Redundant information quantifies only coherent information from multiple sources – for example to have a neuron forward only information that is in agreement between sensory input and internal context information; while unique information reflects what can be obtained from a single source but not from others – for example to make neurons not encode the same information that their peers already encode. Lastly, synergy describes the synthesis of new information from taking multiple sources into account simultaneously – for instance when comparing two signals and computing a difference or error signal. This complexity of operations on input information from multiple classes of sources can be quantified by using the recent framework of “Partial Information Decomposition (PID)” [12, 13, 14]. Given multiple inputs to an output, it can dissect the mutual information into different information atoms, which enumerate the different ways in which the sources can interact to produce the target, as explained above. Through this decomposition, one can paint a comprehensive, yet interpretable picture of how the different input variables contribute to the local computation of an individual neuron.
@@
PID has already been utilized to analyze the information representation and flows in DNNs [15, 16]. We show how this approach can be inverted by using PID to directly formulate goal functions on the individual neuron level. These “infomorphic neurons” then have the ability to optimize for encoding specific parts of the information they receive from their inputs, allowing for an application to tasks from supervised, unsupervised and memory learning, as demonstrated explicitly in [17]. However, neurons with only two different inputs struggle to achieve good performance on classification tasks. It seems that in general, three distinct signals are necessary for local neuron learning: A receptive signal providing the input, a relevance signal helping to filter this input and a distribution signal enabling self-organization between neurons. In classical neural networks, these different classes of information are provided implicitly via the global gradient signals of backpropagation.
@@
The main contributions of this work are as follows: Firstly, we use information theory, and in particular PID, to formulate interpretable, per-neuron goal functions with three input classes, significantly advancing the capabilities of the approach originally presented in [17]. Secondly, we systematically optimize the goal function parameters for these “infomorphic” neurons, provide an intuitive interpretation, and thereby we provide insights into the local computational goals of typical classification tasks. Thirdly, for classification tasks we show that PID-based learning can achieve performance comparable to backpropagation despite being limited to local information only, while being interpretable on a per-neuron basis.
Next, we give a succinct account of infomorphic neurons with two input classes following [17] before generalizing the framework to the more capable version with three input classes in Section 3.
@@ 2 Bivariate infomorphic neurons
Figure 1: Infomorphic neurons are abstract, information-theoretic neurons inspired by the structure of pyramidal neurons [7]. They are trained by adjusting their synaptic weights according to a PID-based goal function.A,B. Inspired by the distinction between apical and basal dendrites in cortical pyramidal neurons, infomorphic neurons are defined as computational units with separate feedforward ($F$) and contextual ($C$) input classes. C: Partial information decomposition (PID) allows one to dissect the total entropy of the neuron into explainable components. D. PID enables one to distinguish how much information comes uniquely from either the feedforward $F$ ($\Pi_{\mathrm{unq,F}}$) or contextual $C$ input ($\Pi_{\mathrm{unq,C}}$) and how much they contribute redundantly ($\Pi_{\mathrm{red}}$), or synergistically ($\Pi_{\mathrm{syn}}$). Classic information theory (top) cannot disentangle these information atoms: The classic entities cover several of the atoms, so that effectively one can only measure redundant minus synergistic information. E: Formulating goal functions $G_{\mathrm{neuron}}$ in terms of PID-atoms $\bm{\Pi}_{\mathrm{neuron}}$ enables one to formulate how strongly redundant, unique, or synergistic information should contribute to a neuron’s output. Figure adapted from [17].
@@
Mutual information can be used to quantify the amount of information that is carried by different input classes about the neuron’s output by considering the aggregated feedforward ($F$), contextual ($C$) and lateral inputs ($L$) as sources $X$ individually. Nevertheless, classical information theory cannot quantify how much of the information is provided redundantly, uniquely or synergistically by different inputs, making it impossible to describe how exactly the information from two source variables contribute to creating the output. Using PID, the total mutual information $I(Y:F,C)$ between the output of the neuron $Y$ and two aggregated inputs, e.g., $F$ and $C$, can be dissected into four PID atoms (see Fig. 1.C): The unique information of the feedforward connection about the output $Y$, $\Pi_{\mathrm{unq,F}}$, is the information which can only be obtained from the feedforward and not the context input, with the unique information $\Pi_{\mathrm{unq,C}}$ of the context being defined analogously. The redundant information $\Pi_{\mathrm{red}}$ reflects the information which can equivalently obtained from either feedforward or contextual inputs about $Y$, while finally the synergistic information $\Pi_{\mathrm{syn}}$ can only be obtained from both inputs considered jointly. All classical mutual information terms between the target and subsets of source variables can be constructed from these PID atoms through the consistency equations [12]
$I(Y:F,C)$ $=\Pi_{\mathrm{red}}+\Pi_{\mathrm{unq,F}}+\Pi_{\mathrm{unq,C}}+\Pi_{\mathrm{syn}}$ (1)
$I(Y:F)$ $=\Pi_{\mathrm{red}}+\Pi_{\mathrm{unq,F}}$
@@
However, these atoms are underdetermined as there are four unknown atoms with only three consistency equations providing constraints [12]. For this reason, an additional quantity needs to be defined, which is usually a measure for redundancy [12]. Such a definition for redundancy, based on the concept of shared exclusions in probability space, is given by the shared-exclusion PID, $I^{\mathrm{sx}}_{\cap}$ [14] (see Section A.2). Importantly, the resulting PID is differentiable with respect to the underlying probability mass function, which allows for a specification and subsequent optimization of learning goals based on PID.
Depending on the task the network is set to solve, different PID atoms become relevant to the information processing. In this work, we focus on the application to supervised learning tasks, in which the ground-truth label is provided as the context $C$ during training. Here, the intuitive goal for the neuron is to foster redundancy between the feedforward and context inputs in its output, to capture only the feedforward signal that agrees with the label. Likewise, if the goal is unsupervised encoding of the input, lateral connections between neurons can be used as the context signal and the goal might become maximizing the unique information of the feedforward input about the output, i.e., to capture only the information which is not already encoded in other neurons [17].
@@
where the residual entropy $H_{\mathrm{res}}=H(Y|F,C)$, is included in the vector $\mathbf{\Pi}$ for brevity of notation. As the $I^{\mathrm{sx}}_{\cap}$ measure is differentiable, the goal function can be optimized by adjusting the weights of the incoming connections of the inputs $F$ and $C$ using gradient ascent – with the analytic formulation of the gradients given in [17], or using a suitable autograd algorithm.
While such infomorphic neurons can be used for supervised learning tasks when each neuron is provided a part of the label to focus on [17], two input classes are insufficient for fully self-organizing the necessary local information processing. This is because if each neuron is tasked with encoding information that is redundant between the whole label and the feedforward signal with no regard to the other neurons, they will inevitably start encoding much of the same information, leading to poor performance. To avoid this, the neurons need to be made aware of what the other neurons encode, similar to the unsupervised example. This can be achieved by incorporating a third input class for lateral connections between neurons of the same layer, as presented next.
@@ 3 Trivariate infomorphic neurons
Figure 2: Adding lateral connections as a third input class enables neurons to self-organize to encode relevant and unique information. A,B. Infomorphic neuron with three inputs - namely feed-forward ($F$), contextual ($C$) and lateral ($L$) inputs. C. With three input classes, the number of PID-atoms increases to 18, represented by different colors, plus the residual entropy $H_{\mathrm{res}}$ in the outer circle. Classical information-theoretic quantities such as the entropy $H(Y)$ and mutual information $I$ with individual sources are depicted by ovals, indicating how they can be built from PID atoms. D. Three input classes allow for the optimization of more complex goals based on 19 different terms (compared to only five in Fig. 1.C). This trivariate PID allows to combine two bivariate objective functions: In a supervised learning task, the intuitive goal is to maximize information in the neuron’s output that is redundant between the feedforward input $F$ and label $C$, while simultaneously ensuring the neuron’s output stays unique with respect to lateral neurons $L$. While bivariate goal functions only allow for optimizing one of these objectives at a time, both objectives can be combined to the goal of maximising the single atom $\Pi_{\{F\}\{C\}}$ in the trivariate case.
Including a third source variable results in a so-called trivariate PID which is more complex than the bivariate case, with atoms in general now describing redundancies between synergies. For example, the atom $\Pi_{\{F\}\{C\}}$ denotes information that can be obtained redundantly from the feedforward and context signal but is unavailable from the lateral inputs, making it ‘uniquely redundant’ to $F$ and $C$. Likewise, the atom $\Pi_{\{F\}\{C,L\}}$ denotes information that can be obtained from $F$ alone but also synergistically from the pair $C,L$ (but not from $C$ or $L$ alone). Overall, upon the introduction of a third source, the four PID atoms of the bivariate setup turn into a total of 18 trivariate PID atoms [13] as shown in Fig. 2.C. These PID atoms $\Pi_{\alpha}$ can in general be addressed by their antichains $\alpha$, i.e., sets of sets of variables, with the inner sets describing synergies and the outer set redundancies between them [13].
Including the residual entropy $H_{\mathrm{res}}$, the decomposition vector $\bm{\Pi}$ and goal parameter vector $\bm{\gamma}$ thus have 19 elements each. The general goal function of a neuron is again given by $G=\bm{\gamma}^{T}\mathbf{\Pi}$ and can be set and optimized in the same way as in the bivariate case:
@@
Utilizing the expressiveness and interpretability of trivariate PID, heuristic goal functions can be derived for different classes of tasks from intuitive reasoning. For the supervised learning task to be solved here, the introduction of the lateral connections as a third input allows to synthesize an intuitive goal function from the two bivariate goals introduced before (see Fig. 2.D). As outlined above, in the bivariate case, canonical goals might either be to foster redundancy between the feedforward and context signals with no regards to what lateral neurons do, or, alternatively, to optimize for uniqueness of the output with respect to the lateral inputs, which does not differentiate between task relevance of information. With trivariate infomorphic neurons, however, these goals can simultaneously be achieved by optimizing for the atom $\Pi_{\{F\}\{C\}}$, which represents the redundancy of the feedforward and context signal which is at the same time unique with respect to the lateral inputs. In simple terms, this enables the neuron to find coherent information between label and input data, that is not yet encoded by other neurons.

### v4

@@ Abstract
In modern deep neural networks, the learning dynamics of individual neurons are often obscure, as the networks are trained via global optimization. Conversely, biological systems build on self-organized, local learning, achieving robustness and efficiency with limited global information. Here, we show how self-organization between individual artificial neurons can be achieved by designing abstract bio-inspired local learning goals. These goals are parameterized using a recent extension of information theory, Partial Information Decomposition (PID), which decomposes the information that a set of information sources holds about an outcome into unique, redundant and synergistic contributions. Our framework enables neurons to locally shape the integration of information from various input classes, i.e., feedforward, feedback, and lateral, by selecting which of the three inputs should contribute uniquely, redundantly or synergistically to the output. This selection is expressed as a weighted sum of PID terms, which, for a given problem, can be directly derived from intuitive reasoning or via numerical optimization, offering a window into understanding task-relevant local information processing. Achieving neuron-level interpretability while enabling strong performance using local learning, our work advances a principled information-theoretic foundation for local learning strategies.
@@ 1 Introduction
Neuroscience models have indeed shown that local learning rules (Hebb, 1949, Dan and Poo, 2004, Song et al., 2000) can solve a variety of tasks, such as sequence learning, unsupervised clustering, and classification tasks (Hebb, 1949, Földiak, 1990, Cramer et al., 2020).
@@
Often formulated on the basis of temporal coincidence (spike-timing), these very mechanistic rules do not offer direct insights into the actual information processing at each neuron. To facilitate such functional insight, more recently information theory has been used to formulate more abstract, but interpretable frameworks for local learning rules (e.g., Kay 2000, Wibral et al. 2017).
Information theory (Shannon, 1948) allows one to abstractly prescribe how much of the information from different inputs should be conveyed to the output of a neuron (Kay, 2000).
@@
Moreover, it provides a general and flexible mathematical framework, which captures the information processing at an abstract and interpretable level free of details of the exact implementation. As such, information theory has been employed to define general optimization goals, including global objectives like cross-entropy loss (Goodfellow et al., 2016) and more localized ones such as local greedy infomax (Löwe et al., 2019), and early attempts at passing only information that is coherent between multiple inputs (Kay and Phillips, 2011).
@@
Nevertheless, classical information theory falls short of describing how multiple sources work together to produce an output. In fact, it completely lacks the capacity to describe how the information from different sources is integrated in redundant, synergistic or unique ways in the output. These different ways, or information atoms, however, contribute very differently to a neuron’s function: Redundant information quantifies only coherent information from multiple sources, for example to have a neuron forward only information that is in agreement between sensory input and internal context information; while unique information reflects what can be obtained from a single source but not from others, for example to make neurons not encode the same information that their peers already encode. Lastly, synergy describes the synthesis of new information from taking multiple sources into account simultaneously, for instance when comparing two signals and computing a difference or error signal.
This complexity of operations on input information from multiple classes of sources can be quantified by using the recent framework of “Partial Information Decomposition (PID)” (Williams and Beer, 2010, Gutknecht et al., 2021, Makkeh et al., 2021). Given multiple inputs to an output, it can dissect the mutual information into different information atoms, which enumerate the different ways in which the sources can interact to produce the target, as explained above. Through this decomposition, one can paint a comprehensive, yet interpretable picture of how the different input variables contribute to the local computation of an individual neuron.
@@
PID has already been utilized to analyze the information representation and flows in DNNs (Ehrlich et al., 2023, Tax et al., 2017). We show how this approach can be inverted by using PID to directly formulate goal functions on the individual neuron level. These “infomorphic neurons” then have the ability to optimize for encoding specific parts of the information they receive from their inputs, allowing for an application to tasks from supervised, unsupervised and memory learning, as demonstrated explicitly in Makkeh et al. (2025). However, neurons with only two different inputs struggle to achieve good performance on classification tasks. It seems that in general, three distinct signals are necessary for local neuron learning: A receptive signal providing the input, a relevance signal helping to filter this input and a distribution signal enabling self-organization between neurons. In classical neural networks, these different classes of information are provided implicitly via the global gradient signals of backpropagation.
@@
The main contributions of this work are as follows: Firstly, we use information theory, and in particular PID, to formulate interpretable, per-neuron goal functions with three input classes, significantly advancing the capabilities of the approach originally presented in Makkeh et al. (2025). Secondly, we systematically optimize the goal function parameters for these “infomorphic” neurons, provide an intuitive interpretation, and thereby obtain insights into the local computational goals of typical classification tasks. Thirdly, for classification tasks we show that PID-based learning can achieve performance comparable to backpropagation despite being limited to local information only, while being interpretable on a per-neuron basis.
Next, we give a succinct account of infomorphic neurons with two input classes following Makkeh et al. (2025) before generalizing the framework to the more capable version with three input classes in Section 3.
@@ 2 Bivariate Infomorphic Neurons
Figure 1: Infomorphic neurons are abstract, information-theoretic neurons inspired by the structure of pyramidal neurons (Wibral et al., 2017). They are trained by adjusting their synaptic weights according to a PID-based goal function.A,B. Inspired by the distinction between apical and basal dendrites in cortical pyramidal neurons, infomorphic neurons are defined as computational units with separate feedforward ($F$) and contextual ($C$) input classes. C. Partial information decomposition (PID) allows one to dissect the total entropy of the neuron into explainable components. D. PID enables one to distinguish how much information comes uniquely from either the feedforward $F$ ($\Pi_{\mathrm{unq,F}}$) or contextual $C$ input ($\Pi_{\mathrm{unq,C}}$) and how much they contribute redundantly ($\Pi_{\mathrm{red}}$), or synergistically ($\Pi_{\mathrm{syn}}$). Classic information theory (top) cannot disentangle these information atoms: The classic entities cover several of the atoms, so that effectively one can only measure redundant minus synergistic information. E. Formulating goal functions $G_{\mathrm{neuron}}$ in terms of PID-atoms $\bm{\Pi}_{\mathrm{neuron}}$ enables one to formulate how strongly redundant, unique, or synergistic information should contribute to a neuron’s output. Figure adapted from Makkeh et al. (2025).
@@
Mutual information can be used to quantify the amount of information that is carried by different input classes about the neuron’s output by considering the aggregated feedforward ($F$), contextual ($C$) and lateral inputs ($L$) as sources $X$ individually. Nevertheless, classical information theory cannot quantify how much of the information is provided redundantly, uniquely or synergistically by different inputs, making it impossible to describe how exactly the information from two source variables contribute to creating the output. Using PID, the total mutual information $I(Y:F,C)$ between the output of the neuron $Y$ and two aggregated inputs, e.g., $F$ and $C$, can be dissected into four PID atoms (see Fig. 1.C): The unique information of the feedforward connection about the output $Y$, $\Pi_{\mathrm{unq,F}}$, is the information which can only be obtained from the feedforward and not the context input, with the unique information $\Pi_{\mathrm{unq,C}}$ of the context being defined analogously. The redundant information $\Pi_{\mathrm{red}}$ reflects the information which can equivalently obtained from either feedforward or contextual inputs about $Y$, while finally the synergistic information $\Pi_{\mathrm{syn}}$ can only be obtained from both inputs considered jointly. All classical mutual information terms between the target and subsets of source variables can be constructed from these PID atoms through the consistency equations (Williams and Beer, 2010)
$I(Y:F,C)$ $=\Pi_{\mathrm{red}}+\Pi_{\mathrm{unq,F}}+\Pi_{\mathrm{unq,C}}+\Pi_{\mathrm{syn}},$ (1)
$I(Y:F)$ $=\Pi_{\mathrm{red}}+\Pi_{\mathrm{unq,F}}\text{ and}$
@@
However, these atoms are underdetermined as there are four unknown atoms with only three consistency equations providing constraints. For this reason, an additional quantity needs to be defined, which is usually a measure for redundancy (Williams and Beer, 2010). Such a definition for redundancy, based on the concept of shared exclusions in probability space, is given by the shared-exclusion PID, $I^{\mathrm{sx}}_{\cap}$ (Makkeh et al., 2021) (see Section A.2). Importantly, the resulting PID is differentiable with respect to the underlying probability mass function, which allows for a specification and subsequent optimization of learning goals based on PID.
Depending on the task the network is set to solve, different PID atoms become relevant to the information processing. In this work, we focus on the application to supervised learning tasks, in which the ground-truth label is provided as the context $C$ during training. Here, the intuitive goal for the neuron is to foster redundancy between the feedforward and context inputs in its output, to capture only the feedforward signal that agrees with the label. Likewise, if the goal is unsupervised encoding of the input, lateral connections between neurons can be used as the context signal and the goal might become maximizing the unique information of the feedforward input about the output, i.e., to capture only the information which is not already encoded in other neurons (Makkeh et al., 2025).
@@
where the residual entropy $H_{\mathrm{res}}=H(Y|F,C)$, is included in the vector $\mathbf{\Pi}$ for brevity of notation. As the $I^{\mathrm{sx}}_{\cap}$ measure is differentiable, the goal function can be optimized by adjusting the weights of the incoming connections of the inputs $F$ and $C$ using gradient ascent—with the analytic formulation of the gradients given in Makkeh et al. (2025), or using a suitable autograd algorithm.
While such infomorphic neurons can be used for supervised learning tasks when each neuron is provided a part of the label to focus on (Makkeh et al., 2025), two input classes are insufficient for fully self-organizing the necessary local information processing. This is because if each neuron is tasked with encoding information that is redundant between the whole label and the feedforward signal with no regard to the other neurons, they will inevitably start encoding much of the same information, leading to poor performance. To avoid this, the neurons need to be made aware of what the other neurons encode, similar to the unsupervised example. This can be achieved by incorporating a third input class for lateral connections between neurons of the same layer, as presented next.
@@ 3 Trivariate Infomorphic Neurons
Figure 2: Adding lateral connections as a third input class enables neurons to self-organize to encode relevant and unique information. A,B. Infomorphic neuron with three inputs - namely feed-forward ($F$), contextual ($C$) and lateral ($L$) inputs. C. With three input classes, the number of PID-atoms increases to 18, represented by different colors, plus the residual entropy $H_{\mathrm{res}}$ in the outer circle. Classical information-theoretic quantities such as the entropy $H(Y)$ and mutual information $I$ with individual sources are depicted by ovals, indicating how they can be built from PID atoms. D. Three input classes allow for the optimization of more complex goals based on 19 different terms (compared to only five in Fig. 1.C). This trivariate PID allows to combine two bivariate objective functions: In a supervised learning task, the intuitive goal is to maximize information in the neuron’s output that is redundant between the feedforward input $F$ and label $C$, while simultaneously ensuring the neuron’s output stays unique with respect to lateral neurons $L$. While bivariate goal functions only allow for optimizing one of these objectives at a time, both objectives can be combined to the goal of maximizing the single atom $\Pi_{\{F\}\{C\}}$ in the trivariate case.
Including a third source variable results in a so-called trivariate PID which is more complex than the bivariate case, with atoms in general now describing redundancies between synergies. For example, the atom $\Pi_{\{F\}\{C\}}$ denotes information that can be obtained redundantly from the feedforward and context signal but is unavailable from the lateral inputs, making it ‘uniquely redundant’ to $F$ and $C$. Likewise, the atom $\Pi_{\{F\}\{C,L\}}$ denotes information that can be obtained from $F$ alone but also synergistically from the pair $C,L$ (but not from $C$ or $L$ alone). Overall, upon the introduction of a third source, the four PID atoms of the bivariate setup turn into a total of 18 trivariate PID atoms (Gutknecht et al., 2021) as shown in Fig. 2.C. These PID atoms $\Pi_{\alpha}$ can in general be addressed by their antichains $\alpha$, i.e., sets of sets of input variables, with the inner sets describing synergies and the outer set redundancies between them (Gutknecht et al., 2021).
Including the residual entropy $H_{\mathrm{res}}=H(Y|F,C,L)$, the decomposition vector $\bm{\Pi}$ and goal parameter vector $\bm{\gamma}$ thus have 19 elements each. The general goal function of a neuron is again given by $G=\bm{\gamma}^{T}\mathbf{\Pi}$ and can be set and optimized in the same way as in the bivariate case:
@@
Utilizing the expressiveness and interpretability of trivariate PID, heuristic goal functions can be derived for different classes of tasks from intuitive reasoning. For the supervised learning task to be solved here, the introduction of the lateral connections as a third input allows to synthesize an intuitive goal function from the two bivariate goals introduced before (see Fig. 2.D). As outlined in the bivariate case above, canonical goals might either be to foster redundancy between the feedforward and context signals with no regards for what lateral neurons do, or, alternatively, to optimize for uniqueness of the output with respect to the lateral inputs, which does not differentiate between task relevance of information. With trivariate infomorphic neurons, however, these goals can simultaneously be achieved by optimizing for the atom $\Pi_{\{F\}\{C\}}$, which represents the redundancy of the feedforward and context signal which is at the same time unique with respect to the lateral inputs. In simple terms, this enables the neuron to find coherent information between label and input data, that is not yet encoded by other neurons.